what would the directrix be for the equation x=y^2-14y+25
Answer by lwsshak3(11628)

what would the directrix be for the equation
x=y^2-14y+25
y^2-14y=x-25
complete the square:
(y^2-14y+49)=x-25+49
(y-7)^2=(x+24)
This is an equation of a parabola that opens right.
Its standard form: (y-k)^2=4p(x-h)
vertex: (-24,7)
axis of symmetry: y=7
4p=1
p=1/4 (distance from vertex to directrix on the axis of symmetry
directrix: x=-97/4